# A Survey on Graph Neural Network Acceleration: Algorithms, Systems, and Customized Hardware

### v1

@@ 2. Notations and Preliminaries
$N$ The total number of nodes in ${\mathcal{G}}$, i.e., $|{\mathcal{V}}|$
@@
2.1. Graph Data Representation
An attributed graph ${\mathcal{G}}$ includes the graph structure $({\mathcal{V}},{\mathcal{E}})$ and node features ${\bm{X}}$, with ${\mathcal{V}}$ denotes a set of nodes, and ${\mathcal{E}}$ denotes a set of edges. A straightforward way to represent the graph structure (edges ${\mathcal{E}}$) is to use an adjacency matrix ${\bm{A}}\in{\mathcal{R}}^{|{\mathcal{V}}|\times|{\mathcal{V}}|}$, with $|{\mathcal{V}}|$ denotes the number of nodes. Many graph algorithms, including the matrix version of GNNs, are presented in terms of ${\bm{A}}$. However, naively representing edges with the full matrix ${\bm{A}}$ requires $O(|{\mathcal{V}}|^{2})$ space complexity, which is impractical to implement for large graphs. In practice, most of the graphs are sparse, i.e., $|{\mathcal{E}}|<<|{\mathcal{V}}|^{2}$. Leveraging the sparsity to represent ${\mathcal{G}}$ in more efficient formats can help save memory and accelerate many computations. One popular sparse representation is the Coordinate (COO) (Wang et al., 2019b) format, which is also known as the “triplet” format. For this format, each edge $(i,j)\in{\mathcal{E}}$ is represented as a triplet $(i,j,w_{ij})$ with $w_{ij}$ being the edge weight for weighted graphs, and $w_{ij}$ equals one for unweighted graphs. The space complexity of a COO graph scales only with the number of edges, e.g., $O(|{\mathcal{E}}|)$. Another popular choice is the Compressed Sparse Row (CSR) format (Saad, 2003; Wang et al., 2019b). It is similar to the COO format but compresses those $|{\mathcal{E}}|$ triplets into three $|{\mathcal{E}}|$-dimensional arrays. One array for all the row indices, e.g., the first entry in COO. One array for all the column indices, e.g., the second entry in COO. One array for all the edge weights, e.g., the third entry in COO. The space complexity of a CSR graph is also $O(|{\mathcal{E}}|)$. Additionally, graphs often come with node features, which are represented as a feature matrix ${\bm{X}}\in{\mathcal{R}}^{|{\mathcal{V}}|\times D}$. ${\bm{X}}$ is often referred to as the raw features to distinguish them from the hidden features learned by the model. Similarly, there could be features for graph edges, but we do not assume edge features are given in this survey because they are not generally available.
2.2. Graph Neural Networks
@@
where the matrix ${\bm{H}}^{(l-1)}$ contains hidden representations of all nodes at the $(l-1)$-th step. It is first multiplied by a filter matrix ${\bm{P}}$ (on the left) and a neural network weight matrix ${\bm{W}}^{(l)}$ (on the right) and then is passed through a nonlinear activation function $\sigma(\cdot)$. ${\bm{P}}$ corresponds to the AGGR function and ${\bm{W}}^{(l)}$ and $\sigma(\cdot)$ correspond to the UPDATE function (when it is an MLP) in Equation (1). ${\bm{P}}$ can take different forms for different GNNs, and it is usually a function of ${\bm{A}}$. For example, ${\bm{P}}$ is a degree-normalized version of the adjacency matrix with self-loops for GCN, i.e., ${\bm{P}}=\tilde{\bm{D}}^{-\frac{1}{2}}\tilde{\bm{A}}\tilde{\bm{D}}^{-\frac{1}{2}}$.
@@
The matrix version looks cleaner but it is implemented less often in practice since the full matrix representation of ${\bm{P}}$ usually has space complexity $O(|{\mathcal{V}}|^{2})$, so it is infeasible to fit such ${\bm{P}}$ in memory for large graphs. More frequently, ${\bm{P}}$ is represented as a sparse matrix (e.g., in COO or CSR format) mentioned above, and message passing is implemented using the Scatter-ApplyEdge-Gather-ApplyVertex (SAGA) paradigm (Ma et al., 2019a). For SAGA, messages are first computed and stored on edges, and then each node aggregates messages from its edges and updates the representation. We show the SAGA GNN message passing in Algorithm 1, which has time complexity $O(|{\mathcal{E}}|)$.
2.3. Graph Machine Learning Tasks
@@
3. GNN Acceleration: Training Algorithms
In this section, we discuss two categories of training acceleration techniques: graph modification and sampling. Both methods aim to reduce the computation graph to accelerate training. The difference is whether there is a modified graph ${\mathcal{G}}^{\prime}$ as an intermediate output. For graph modification, the procedure can be separated into two steps. The first step outputs a graph ${\mathcal{G}}^{\prime}$ that is smaller than ${\mathcal{G}}$ and can be fastly trained on. The second stage is regular GNN training with ${\mathcal{G}}^{\prime}$ as the input. For sampling, a subset of nodes/edges is selected to construct smaller computation graphs in each training iteration. Generally speaking, sampling also modifies ${\mathcal{G}}$, but the modification is done dynamically and implicitly, so there is no intermediate ${\mathcal{G}}^{\prime}$ output. Since the computation graph can differ from iteration to iteration, all the nodes and edges have the chance to be covered by sampling. In contrast, for graph modification, not all the nodes and edges will appear in ${\mathcal{G}}^{\prime}$, but new nodes and edges may be created.
3.1. Graph Modification
Graph modification accelerates GNN training in two steps. The first step takes the graph ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$, node features ${\bm{X}}$, and labels ${\bm{Y}}$ as input and outputs modified ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime})$, ${\bm{X}}^{\prime}$, and ${\bm{Y}}^{\prime}$ for training. Training on ${\mathcal{G}}^{\prime}$ will be faster, if ${\mathcal{G}}^{\prime}$ is smaller than ${\mathcal{G}}$ in the sense that $|{\mathcal{V}}^{\prime}|<|{\mathcal{V}}|$ and/or $|{\mathcal{E}}^{\prime}|<|{\mathcal{E}}|$. The second step trains a GNN on ${\mathcal{G}}^{\prime}$ such that the GNN has a similar performance on ${\mathcal{G}}$ as if it is trained on ${\mathcal{G}}$. Mathematically, regular training outputs $\theta^{*}=\operatorname*{arg\,min}_{\theta}{\mathcal{L}}(\textsc{GNN}_{\theta}({\mathcal{G}},{\bm{X}}),{\bm{Y}})$, and training with graph modification outputs $\theta^{\prime*}=\operatorname*{arg\,min}_{\theta^{\prime}}{\mathcal{L}}(\textsc{GNN}_{\theta^{\prime}}({\mathcal{G}}^{\prime},{\bm{X}}^{\prime}),{\bm{Y}}^{\prime})$ such that $\textsc{GNN}_{\theta^{\prime*}}$ has similar performance as $\textsc{GNN}_{\theta^{*}}$ on ${\mathcal{G}}$. In the following, we go over several types of graph modification methods including graph coarsening, graph sparsification, and graph condensation. Each of these methods modifies the graph ${\mathcal{G}}$ differently to create its own ${\mathcal{G}}^{\prime}$, but all ${\mathcal{G}}^{\prime}$ are smaller graphs that accelerate GNN training. An illustration of these graph modification methods is shown in Figure 2.
Figure 2. Illustration of graph modification methods: Graph Coarsening methods perform graph clustering and merge clusters of nodes into a super-node. Graph Sparsification methods remove less important edges. Graph Condensation methods generate a new condensed graph using a randomly initialized generative model. For the modified graph (the rightmost column), black nodes/edges are from the original graph, and colored nodes/edges are newly created.
3.1.1. Graph Coarsening
Graph coarsening is a technique aiming at reducing the graph size while preserving its overall structure. Given a graph ${\mathcal{G}}$, graph coarsening merges nodes in the same local structure of ${\mathcal{G}}$ into a single “super-node” and merges edges connecting super-nodes into a “super-edge”. The coarse graph ${\mathcal{G}}^{\prime}$ thus has much fewer nodes and edges compared to ${\mathcal{G}}$. The key step of coarsening is to cluster nodes in ${\mathcal{G}}$ into $K$ clusters for merging. Nodes in the same cluster will be merged into a super-node, and then a super-edge $(k,o)$ connecting super-node $k$ and super-node $o$ will be constructed if there are edges connecting nodes in cluster $k$ and nodes in cluster $o$. An illustration of the clustering and merging of graph coarsening is shown in Figure 2 top row.
Graph coarsening, with its central step being graph clustering, has been studied long before GNNs were proposed (Kernighan and Lin, 1970; Karypis and Kumar, 1998; Von Luxburg, 2007), and new coarsening algorithms have been proposed until very recently (Loukas and Vandergheynst, 2018; Loukas, 2019; Bravo Hermsdorff and Gunderson, 2019). The key idea of these algorithms is to coarsen the graph while preserving some graph properties, which are often related to the graph spectrum, e.g., restricted spectral approximation (Loukas and Vandergheynst, 2018) and inverse Laplacian (Bravo Hermsdorff and Gunderson, 2019). Below, we discuss the graph coarsening algorithms related to GNNs.
@@ 3.1.1. Graph Coarsening
Huang et al. (Huang et al., 2021b) discuss a general framework using graph coarsening to accelerate GNN training. In particular, a graph ${\mathcal{G}}$ is partitioned into K clusters with sizes $N_{1},...,N_{K}$. A matrix ${\bm{C}}\in{\mathcal{R}}^{N\times K}$ is used to represent the partition, where ${\bm{C}}_{i,j}=1/\sqrt{N_{j}}$ if and only if node $i$ is assigned to cluster $j$, and ${\bm{C}}_{i,j}=0$ otherwise, so that ${\bm{C}}^{\top}{\bm{C}}={\bm{I}}$. The super-nodes and super-edges are constructed according to ${\bm{C}}$. Moreover, the super-node features ${\bm{X}}^{\prime}$ are set to be the weighted average of node features within each cluster, i.e., ${\bm{X}}^{\prime}={\bm{C}}^{\top}{\bm{X}}$. The super-nodes labels ${\bm{Y}}^{\prime}$ are constructed by taking the dominant label of nodes in each super-node, i.e., ${\bm{y}}^{\prime}_{k}=\textsc{ONEHOT}(\operatorname*{arg\,max}(({\bm{C}}^{\top}{\bm{Y}})_{k}))$. (Huang et al., 2021b) considers different partition algorithms for coarsening the graph: spectral clustering (Von Luxburg, 2007) and several partition algorithms from (Loukas, 2019) with restricted spectral approximation. Empirically, the Variation Neighborhoods method from (Loukas, 2019) works better than other partition algorithms.
There are some other related works on the topic of graph coarsening and GNNs, but their goals are different from our focus on graph coarsening for GNN training acceleration. GOREN (Cai et al., 2021) uses GNNs as a tool to help do better graph coarsening, where the goal is to output a coarse graph with better Rayleigh loss and Eigenerror. GraphZoom (Deng et al., 2020) uses graph coarsening to improve the accuracy and scalability of graph embedding algorithms like DeepWalK (Perozzi et al., 2014) and Node2Vec (Grover and Leskovec, 2016).
3.1.2. Graph Sparsification
Graph sparsification by removing redundant edges is a natural idea of graph modification to reduce the computation graph and accelerate GNN training. An illustration of graph sparsification is shown in Figure 2 middle row. Like graph coarsening, graph sparsification has been studied and widely used before GNNs became popular (Benczúr and Karger, 1996; Spielman and Srivastava, 2011; Spielman and Teng, 2014; Serrano et al., 2009). Graph sparsification methods usually try to find a sparsified ${\mathcal{G}}^{\prime}$ that can well approximate the properties of the original ${\mathcal{G}}$. That is keeping all the nodes in ${\mathcal{G}}$ and removing less-important edges without disconnecting ${\mathcal{G}}$. Example properties preserved by existing graph sparsification methods include the total weight of cuts (Benczúr and Karger, 1996), spectral properties (Spielman and Srivastava, 2011; Spielman and Teng, 2014), and hierarchical structures (Serrano et al., 2009).
With the goal being GNN training acceleration, the above graph sparsification algorithms can be applied as a pre-processing step before GNN training starts. Since the sparsification algorithms are usually highly scalable and only need to be run once, the time it takes is negligible compared to GNN training. A sparsified graph can reduce the full-batch time complexity of training an $L$-layer GNN from $O(L|{\mathcal{E}}|)$ to $O(L|{\mathcal{E}}^{\prime}|)$, where $|{\mathcal{E}}^{\prime}|$ could be much smaller than $|{\mathcal{E}}|$.
There are some related works performing GNN-based graph sparsification to improve GNN accuracy (Li et al., 2020) or robustness (Zheng et al., 2020b) but not for GNN acceleration explicitly. Their difference from the traditional methods is the sparsification criteria consider both ${\mathcal{G}}$ and the GNN instead of the properties of ${\mathcal{G}}$ only. These criteria are usually formed as optimization problems and solved while training the GNN. For example, (Li et al., 2020) finds a sparsified graph ${\mathcal{G}}^{\prime}$ via the alternating direction method of multipliers. (Zheng et al., 2020b) learns a sparsification strategy by generating sparse k-neighbor subgraphs. It is not guaranteed that these sparsification methods will always accelerate single GNN training. Although the sparse graph reduces ${\mathcal{E}}$ to accelerate message passing and is likely to make the model training converge faster, solving the extra optimization problem takes time and can slow down the training. These methods may be used to accelerate the training of the GNNs after the first one if multiple GNNs need to be trained on the same graph, as for the neural architecture search task.
A closely related concept that may be confused with sparsification is pruning. Some papers use these two terms interchangeably, but conventionally pruning refers to a modification of the neural network model in the deep learning context. Thus, model pruning and graph pruning are totally different concepts with pruning more often referring to the former. We will discuss model pruning in more detail in Section 4.1 for GNN inference acceleration. For graph pruning, generally speaking, it is broader than graph sparsification and includes pruning graph nodes as well. However, since GNN is trained to learn node representations, it is rare to prune nodes in ${\mathcal{G}}$ completely without preserving their information. What is more often done is dynamically exclude different nodes in each training iteration, but each node has a probability to be included in the entire training process. This is the sampling procedure that we will discuss in Section 3.2.
3.1.3. Graph Condensation
Graph condensation generates a completely new graph that preserves the training dynamics of the original graph (Jin et al., 2021, 2022). In other words, the result of training two GNNs with identical architecture on the original graph ${\mathcal{G}}$ and the condensed graph ${\mathcal{G}}^{\prime}$ (the generated new graph) respectively should match. Meanwhile, ${\mathcal{G}}^{\prime}$ can be much smaller than ${\mathcal{G}}$ and thus accelerating GNN training. An illustration of graph condensation is shown in Figure 2 bottom row.
GCond (Jin et al., 2021) matches the training gradients of two GNNs to generate a condensed graph. In particular, a randomly initialized generative model first generates the condensed graph ${\mathcal{G}}^{\prime}=({\mathcal{V}}^{\prime},{\mathcal{E}}^{\prime},{\bm{X}}^{\prime})$ and labels ${\bm{Y}}^{\prime}$, given a condensation ratio $r$, and two GNNs $GNN_{\theta}$ and $GNN_{\theta^{\prime}}$ are trained in parallel using stochastic gradient descent (SGD) on ${\mathcal{G}}$ and ${\mathcal{G}}^{\prime}$ respectively. At each training step, SGD generates gradients $\nabla_{\theta}{\mathcal{L}}(GNN_{\theta}({\mathcal{G}}),{\bm{Y}})$ and $\nabla_{\theta^{\prime}}{\mathcal{L}}(GNN_{\theta^{\prime}}({\mathcal{G}}^{\prime}),{\bm{Y}}^{\prime})$ for updating $\theta$ and $\theta^{\prime}$ respectively. Then gradient matching is performed by tuning the generative model to minimize the distance between $\nabla_{\theta}{\mathcal{L}}(GNN_{\theta}({\mathcal{G}}),{\bm{Y}})$ and $\nabla_{\theta^{\prime}}{\mathcal{L}}(GNN_{\theta^{\prime}}({\mathcal{G}}^{\prime}),{\bm{Y}}^{\prime})$, so that ${\mathcal{G}}^{\prime}$ and ${\bm{Y}}^{\prime}$ better preserve the training dynamics can be generated in the next iteration. For the generative model, since ${\mathcal{E}}^{\prime}$ and ${\bm{X}}^{\prime}$ depend on each other in general graphs, GCond chooses to generate ${\bm{X}}^{\prime}$ first and then construct ${\mathcal{E}}^{\prime}$ from ${\bm{X}}^{\prime}$. Specifically, GCond initializes $r|{\mathcal{V}}|$ node features ${\bm{X}}^{\prime}$ as free learnable parameters, which will be directly updated by the gradient-matching loss. To generate condensed edges ${\mathcal{E}}^{\prime}$, GCond uses an MLP as a feature-based structure prediction model, which takes input ${\bm{x}}^{\prime}_{i}$ and ${\bm{x}}^{\prime}_{j}$ and predicts edge $i$-$j$. Finally, for condensed labels ${\bm{Y}}^{\prime}$, GCond initializes a ${\bm{y}}^{\prime}$ for each ${\bm{x}}^{\prime}_{i}$ and fixes ${\bm{y}}^{\prime}$ by randomly selecting a class following the label distribution in the original graph ${\mathcal{G}}$ without learning. Since all ${\bm{X}}^{\prime}$ are randomly initialized, it doesn’t matter what is the value of ${\bm{y}}^{\prime}$ assigned to each ${\bm{x}}^{\prime}_{i}$, and only the distribution of ${\bm{Y}}^{\prime}$ matters. GCond sets the distribution ${\bm{Y}}^{\prime}$ to follow the original ${\bm{Y}}$ so that the class ratios in ${\mathcal{G}}$ are kept in ${\mathcal{G}}^{\prime}$.
3.2. Sampling
Sampling is a prevalent technique for accelerating GNN training, especially under computation resource limitation (Liu et al., 2021b; Serafini and Guan, 2021). The basic idea is to dynamically select a subset of nodes and edges during each training iteration and build the computation graph only with the selected nodes and edges. Sampling can greatly reduce the computation graph size and accelerate model training. It is found that the model accuracy can also be maintained by sampling since the full neighborhood is not always required to predict a target node. The shared goal of different sampling algorithms is to reduce the time and memory consumption as much as possible while maintaining model accuracy. Note that sampling is different from the graph modification discussed above. Sampling is done dynamically and implicitly and there is no intermediate modified graph output.
In the following, we introduce different sampling algorithms in a unified framework. Throughout the section, we use $\textsc{SAMPLE}(\cdot)$ to denote performing the sampling operation and use $\overline{\cdot}$ on our notations to denote the sampled version. Given each target node $v$, instead of aggregating from all its full neighborhood $\mathcal{N}(v)$ for conducting message passing, sampling selects a subset of neighboring nodes, denoted $\overline{\mathcal{N}}(v)$, which provides the most important information for predicting $v$. A sampled version of GNN message passing in Equation (1) and Equation (2) can approximate the node representations as:
(4) ${\bm{h}}_{v}^{(l)}$ $\approx\textsc{UPDATE}({\bm{h}}_{v}^{(l-1)},\textsc{AGGR}(\{\phi({\bm{h}}_{v}^{(l-1)},{\bm{h}}_{u}^{(l-1)})|u\in\overline{{\mathcal{N}}}(v)^{(l)}\}))$
(5) ${\bm{H}}^{(l)}$ $\approx\sigma(\overline{{\bm{P}}}^{(l)}{\bm{H}}^{(l-1)}{\bm{W}}^{(l)})$
where $\overline{{\mathcal{N}}}(v)^{(l)}$ and $\overline{{\bm{P}}}^{(l)}$ indicate the layer-$l$-sampled neighborhood and its corresponding filter matrix respectively.
The central idea of sampling is to find informative neighborhoods that can construct small but effective computation graphs. Existing sampling algorithms can be categorized into three types according to their input to the $\textsc{SAMPLE}(\cdot)$ operation: 1) node-wise, which takes a single node neighborhood as input and sample within it; 2) layer-wise, which takes all nodes in the same computation graph layer and sample all their neighborhoods jointly; 3) subgraph-wise, which takes a whole graph as input and sample a subgraph as neighborhoods for nodes in the graph. Figure 3 illustrates the workflow of these three types of sampling algorithms. In the following, we discuss the details of sampling algorithms in each category and summarize their pros and cons.

### v2

@@ Appendix A Notations and Preliminaries
NN The total number of nodes in $\mathcal{G}$}, i.e., $|{\mathcal{V}}|$
@@
A.1. Graph Data Representation
An attributed graph $\mathcal{G}$} includes the graph structure $({\mathcal{V}},{\mathcal{E}})$ and node features $\bm{X}$}, with $\mathcal{V}$} denotes a set of nodes, and $\mathcal{E}$} denotes a set of edges. A straightforward way to represent the graph structure (edges $\mathcal{E}$}) is to use an adjacency matrix ${\bm{A}}\in{\mathcal{R}}^{|{\mathcal{V}}|\times|{\mathcal{V}}|}$, with $|{\mathcal{V}}|$ denotes the number of nodes. Many graph algorithms, including the matrix version of GNNs, are presented in terms of $\bm{A}$}. However, naively representing edges with the full matrix $\bm{A}$} requires $O(|{\mathcal{V}}|^{2})$ space complexity, which is impractical to implement for large graphs. In practice, most of the graphs are sparse, i.e., $|{\mathcal{E}}|<<|{\mathcal{V}}|^{2}$. Leveraging the sparsity to represent $\mathcal{G}$} in more efficient formats can help save memory and accelerate many computations. One popular sparse representation is the Coordinate (COO) (Wang et al., 2019a) format, which is also known as the “triplet” format. For this format, each edge $(i,j)\in{\mathcal{E}}$ is represented as a triplet $(i,j,w_{ij})$ with $w_{ij}$ being the edge weight for weighted graphs, and $w_{ij}$ equals one for unweighted graphs. The space complexity of a COO graph scales only with the number of edges, e.g., $O(|{\mathcal{E}}|)$. Another popular choice is the Compressed Sparse Row (CSR) format (Saad, 2003; Wang et al., 2019a). It is similar to the COO format but compresses those $|{\mathcal{E}}|$ triplets into three $|{\mathcal{E}}|$-dimensional arrays. One array for all the row indices, e.g., the first entry in COO. One array for all the column indices, e.g., the second entry in COO. One array for all the edge weights, e.g., the third entry in COO. The space complexity of a CSR graph is also $O(|{\mathcal{E}}|)$. Additionally, graphs often come with node features, which are represented as a feature matrix ${\bm{X}}\in{\mathcal{R}}^{|{\mathcal{V}}|\times D}$. $\bm{X}$} is often referred to as the raw features to distinguish them from the hidden features learned by the model. Similarly, there could be features for graph edges, but we do not assume edge features are given in this survey because they are not generally available.
A.2. Graph Neural Networks
@@
where the matrix ${\bm{H}}^{(l-1)}$ contains hidden representations of all nodes at the (l1)(l-1)-th step. It is first multiplied by a filter matrix $\bm{P}$} (on the left) and a neural network weight matrix ${\bm{W}}^{(l)}$ (on the right) and then is passed through a nonlinear activation function $\sigma(\cdot)$. $\bm{P}$} corresponds to the AGGR function and ${\bm{W}}^{(l)}$ and $\sigma(\cdot)$ correspond to the UPDATE function (when it is an MLP) in Equation (8). $\bm{P}$} can take different forms for different GNNs, and it is usually a function of $\bm{A}$}. For example, $\bm{P}$} is a degree-normalized version of the adjacency matrix with self-loops for GCN, i.e., ${\bm{P}}=\tilde{\bm{D}}^{-\frac{1}{2}}\tilde{\bm{A}}\tilde{\bm{D}}^{-\frac{1}{2}}$.
@@
The matrix version looks cleaner but it is implemented less often in practice since the full matrix representation of $\bm{P}$} usually has space complexity $O(|{\mathcal{V}}|^{2})$, so it is infeasible to fit such $\bm{P}$} in memory for large graphs. More frequently, $\bm{P}$} is represented as a sparse matrix (e.g., in COO or CSR format) mentioned above, and message passing is implemented using the Scatter-ApplyEdge-Gather-ApplyVertex (SAGA) paradigm (Ma et al., 2019b). For SAGA, messages are first computed and stored on edges, and then each node aggregates messages from its edges and updates the representation. We show the SAGA GNN message passing in Algorithm 1, which has time complexity $O(|{\mathcal{E}}|)$.
A.3. Graph Machine Learning Tasks
@@
Appendix B Detailed Discussion of Training Algorithms
B.1. Graph modification
@@
Huang et al. (Huang et al., 2021a) discuss a general framework using graph coarsening to accelerate GNN training. In particular, a graph $\mathcal{G}$} is partitioned into K clusters with sizes $N_{1},...,N_{K}$. A matrix ${\bm{C}}\in{\mathcal{R}}^{N\times K}$ is used to represent the partition, where ${\bm{C}}_{i,j}=1/\sqrt{N_{j}}$ if and only if node $i$ is assigned to cluster $j$, and ${\bm{C}}_{i,j}=0$ otherwise, so that ${\bm{C}}^{\top}{\bm{C}}={\bm{I}}$. The super-nodes and super-edges are constructed according to $\bm{C}$}. Moreover, the super-node features $\bm{X}$$^{\prime}$ are set to be the weighted average of node features within each cluster, i.e., ${\bm{X}}^{\prime}={\bm{C}}^{\top}{\bm{X}}$. The super-nodes labels $\bm{Y}$$^{\prime}$ are constructed by taking the dominant label of nodes in each super-node, i.e., ${\bm{y}}^{\prime}_{k}=\textsc{ONEHOT}(\operatorname*{arg\,max}(({\bm{C}}^{\top}{\bm{Y}})_{k}))$. (Huang et al., 2021a) considers different partition algorithms for coarsening the graph: spectral clustering (Von Luxburg, 2007) and several partition algorithms from (Loukas, 2019) with restricted spectral approximation. Empirically, the Variation Neighborhoods method from (Loukas, 2019) works better than other partition algorithms.
B.2. Sampling
B.2.1. Benchmark Datasets